# Training Nonlinear Transformers for Chain-of-Thought Inference: A Theoretical Generalization Analysis

### v2

@@ 1 Introduction
Transformer-based large-scale foundation models, such as GPT-3 (Brown et al., 2020), GPT-4 (OpenAI, 2023), LLaMa (Touvron et al., 2023a, b), and Sora (Liu et al., 2024), have demonstrated remarkable success across various tasks, including natural language processing (Brown et al., 2020; Touvron et al., 2023b), multimodal learning (OpenAI, 2023; Radford et al., 2021), and image/video generation (OpenAI, 2023; Liu et al., 2024). What is more surprising is that large language models (LLMs) demonstrate reasoning ability through the so-called “Chain-of-Thought” (CoT) method (Wei et al., 2022). The objective is to let a pre-trained LLM generate $K$ steps of reasoning given input query ${\bm{x}}_{query}$ without any fine-tuning. To achieve that, the input ${\bm{x}}_{query}$ is augumented with $l$ examples $\{{\bm{x}}_{i},\{{\bm{y}}_{i,j}\}_{j=1}^{K}\}_{i=1}^{l}$ of a certain $K$-step reasoning task, where each ${\bm{x}}_{i}$ is the input with ${\bm{y}}_{i,j}$ as the $j$-th reasoning step, and ${\bm{y}}_{i,K}$ is the final output. A pre-trained model then takes the resulting augmented input, referred to as a prompt, and outputs the corresponding reasoning steps $\{{\bm{z}}_{j}\}_{j=1}^{K}$ for ${\bm{x}}_{query}$, or simply outputs ${\bm{z}}_{K}$. CoT can be viewed as an extended and more intelligent method than the previous in-context learning (ICL) method, where only input-label pairs $\{{\bm{x}}_{i},{\bm{y}}_{i,K}\}_{i=1}^{l}$ are augmented in the prompt to predict ${\bm{z}}_{K}$ with the pre-trained model.
Inspired by the outstanding empirical performance of CoT in arithmetic reasoning (Wang et al., 2023; Zhang et al., 2023b; Wang and Zhou, 2024), symbolic reasoning (Zhang et al., 2023b; Zhou et al., 2023), and commonsense reasoning (Wang et al., 2023; Wang and Zhou, 2024), there have been some recent works (Li et al., 2023d; Feng et al., 2023; Li et al., 2024d; Yang et al., 2024; Wen et al., 2024) on the theoretical understanding of CoT. These works investigate CoT from the perspective of expressive power, i.e., they construct the Transformer architecture that is proven to have the CoT ability. They also demonstrate empirically that supervised training on pairs of CoT prompts and corresponding outputs can lead to models with CoT ability. However, none of these results theoretically address the question of why a Transformer can obtain generalization-guaranteed CoT ability by training from data with gradient-based methods. Meanwhile, another line of research (Zhang et al., 2023a; Huang et al., 2023; Wu et al., 2023; Li et al., 2024a) aims to unveil the reasons behind the ICL ability of Transformers through characterizing the training dynamics of a Transformer in the supervised setting. These analyses are specifically applicable to ICL. Therefore, a theoretical question still remains less explored, i.e.,
@@ 1.1 Major Contributions
Following Li et al. (2023d); Feng et al. (2023); Li et al. (2024d); Yang et al. (2024); Wen et al. (2024), we train the model in a supervised setting using prompt and label pairs. This paper provides the first theoretical analysis of the training dynamics of nonlinear Transformers to achieve CoT ability. We prove that the learned model has guaranteed CoT ability for new tasks with distribution shifts from the training tasks, even when there exist noisy and erroneous context examples in the prompt. We theoretically characterize the required number of training samples and iterations needed to train a desirable model and the number of context examples required for successful CoT reasoning with a generalization guarantee. Moreover, we provide a theoretical explanation for why CoT outperforms ICL in some cases. Our main technical contributions are as follows:
@@ 1.2 Related works
Expressive power of CoT Li et al. (2023d) proves the existence of a Transformer that can learn a multi-layer perceptron (MLP). They interpret CoT as first filtering important tokens and then making predictions by ICL. They also establish the required number of context examples for a desired prediction with the constructed Transformer. Feng et al. (2023); Li et al. (2024d); Merrill and Sabharwal (2024) show that Transformers with CoT are more expressive than Transformers without CoT. Yang et al. (2024); Wen et al. (2024) show the superiority of standard Transformers in some reasoning tasks compared with recurrent neural networks and linear Transformers.
@@
Training and Generalization of Transformers There have been several recent works about the optimization and generalization analysis of Transformers. Jelassi et al. (2022); Li et al. (2023c); Oymak et al. (2023); Li et al. (2023a, b, 2024b); Luo (2023); Huang et al. (2024); Zhang et al. (2024) study the generalization of one-layer Transformers by assuming spatial association, semantic/contextual structure, or the majority voting of tokens in the data. Oymak et al. (2023); Tarzanagh et al. (2023b, a); Tian et al. (2023a, b); Li et al. (2024c); Ildiz et al. (2024); Nichani et al. (2024); Makkuva et al. (2024b) investigate the training dynamics or loss landscape of Transformers for the next token prediction by assuming infinitely long input sequences, causal structure/Markov Chain of data, or a proper prediction head. Deora et al. (2023); Chen and Li (2024) analyze the optimization and generalization of multi-head attention networks.
@@ 2.1 Training to acquire the Chain-of-Thought ability
The learning model is a single-head, one-layer attention-only Transformer. We consider positional encoding $\{{\bm{c}}_{k}\}_{k=1}^{K}\in\mathbb{R}^{2d_{\mathcal{X}}}$. Following theoretical works (Jelassi et al., 2022; Huang et al., 2024; Ildiz et al., 2024), we add the positional encoding to each ${\bm{p}}_{i}$ by $\tilde{{\bm{p}}}_{i}={\bm{p}}_{i}+{\bm{c}}_{(i\mod K)}$, $i\in[K(l_{tr}+1)]$. $\tilde{{\bm{p}}}_{query}$ is also defined by adding the corresponding ${\bm{c}}_{k}$ to ${\bm{p}}_{query}$. Mathematically, given a prompt ${\bm{P}}$ defined in (1) with $\text{len}(P)$ (which is at most $K(l_{tr}+1)$) denoting the number of columns, it can be written as
$F(\Psi;{\bm{P}})=\sum_{i=1}^{\text{len}(P)-1}{\bm{W}}_{V}\tilde{{\bm{p}}}_{i}\cdot\text{softmax}(({\bm{W}}_{K}\tilde{{\bm{p}}}_{i})^{\top}{\bm{W}}_{Q}\tilde{{\bm{p}}}_{query}),$ (2)
where ${\bm{W}}_{Q},{\bm{W}}_{K}\in\mathbb{R}^{m\times(2d_{\mathcal{X}})}$, ${\bm{W}}_{V}\in\mathbb{R}^{d_{\mathcal{X}}\times(2d_{\mathcal{X}})}$ are the embedding matrices for queries, keys, and values, respectively. $\Psi:=\{{\bm{W}}_{Q}$, ${\bm{W}}_{K},{\bm{W}}_{V}\}$ is the set of all model weights [^1] We focus on a one-layer single-head Transformer motivated by recent advancements and current state in Transformer and CoT analysis. Please see Appendix A.1 for discussion. . Typically, $m>2d_{\mathcal{X}}$.
@@ 2.2 Training Algorithm
For simplicity of analysis, we let ${\bm{W}}={\bm{W}}_{K}^{\top}{\bm{W}}_{Q}$ and ${\bm{W}}_{V}=(\bm{0}_{d_{\mathcal{X}}\times d_{\mathcal{X}}}\ {\bm{I}}_{d_{\mathcal{X}}})\in\mathbb{R}^{d_{\mathcal{X}}\times(2d_{\mathcal{X}})}$ as (Jelassi et al., 2022; Huang et al., 2023; Zhang et al., 2023a; Huang et al., 2024). Let $\{{\bm{c}}_{k}\}_{k=1}^{K}$ be a set of orthonormal vectors. The model is trained using stochastic gradient descent (SGD) with step size $\eta$ with batch size $B$, summarized in Algorithm 1 in Appendix B. Each entry of ${\bm{W}}^{(0)}$ is generated from $\mathcal{N}(0,\xi^{2})$ for a tiny $\xi>0$. ${\bm{W}}_{V}$ is fixed during the training. The fraction of prompts with ${\bm{z}}_{k-1}$ as the query input is $1/K$ by uniform sampling for any $k\in[K]$ in each batch.
@@ 2.3 Chain-of-Thought Inference
We follow the CoT-I/O scheme formulated in (Li et al., 2023d; Feng et al., 2023; Li et al., 2024d; Yang et al., 2024; Park et al., 2024) as the inference method. Specifically, for a $K$-step CoT with $l_{ts}$ examples on a certain $f\in\mathcal{T}^{\prime}$, given the testing prompt ${\bm{P}}$ defined in (4), let ${\bm{P}}_{1}={\bm{P}}$ and ${\bm{P}}_{0}$ be the first $K\cdot l_{ts}$ columns of ${\bm{P}}$. When we use CoT prompting for prediction in the $k$-th step, we first generate the output ${\bm{v}}_{k}$, $k\in[K]$ via greedy decoding by feeding the $k$-th step prompt ${\bm{P}}_{k}$ to the trained model $\Psi$ obtained from (3). The greedy decoding scheme means outputting the most probable token from the discrete set $\mathcal{Y}$ of all possible outputs, as stated in (5).
@@
where ${\bm{q}}_{k}$ is the $k$-th step reasoning column for the query. The model finally outputs ${\bm{v}}_{1},\cdots,{\bm{v}}_{K}$ as CoT result for query ${\bm{x}}_{query}$ by (5). The CoT process is summarized in Algorithm 2 of Appendix B.
When $K\geq 2$, following (Li et al., 2023d; Feng et al., 2023; Li et al., 2024d; Yang et al., 2024), the CoT generalization error given the testing query ${\bm{x}}_{query}$, the testing data distribution $\mathcal{D}^{\prime}$, and the labels $\{{\bm{z}}_{k}\}_{k=1}^{K}$ on a $K$-steps testing task $f\in\mathcal{T}^{\prime}$ is defined as
@@ 3.1 Main Theoretical Insights
We consider the setup that the model is trained using samples generated from tasks in $\mathcal{T}$ that operate on $M$ orthonormal training-relevant (TRR) patterns, while both CoT and ICL are evaluated on tasks in $\mathcal{T}^{\prime}$ that operate on $M^{\prime}$ orthonormal testing-relevant (TSR) patterns that belong to the span of TRR patterns. We consider the general setup that the context examples in the prompt for CoT and ICL testing are both noisy, i.e., TSR patterns with additive noise, and partially inaccurate, i.e., the reasoning in some examples contains incorrect steps. Our main insights are as follows.
@@
P2. Guaranteed CoT Generalization. To achieve zero CoT error on tasks in $\mathcal{T}^{\prime}$ with the learned model, Theorem 2 shows that the required number of context examples, where noise and errors are present, for task $f$ in the testing prompt is proportional to ${(\alpha^{\prime}\tau^{f}\rho^{f})}^{-2}$, where $\alpha^{\prime}$ is the fraction of context examples with inputs that share the same TSR patterns as the query, the constant $\tau^{f}$ in $(0,1)$ measures the fraction of accurate context examples, and a larger constant $\rho^{f}$ in $(0,1)$ reflects a higher reasoning accuracy in each step of the examples. This result formally characterizes the intuition that more accurate context examples and more similar examples to the query improve the CoT accuracy.
@@ 3.2 The Formulation of Data and Tasks
Every training prompt ${\bm{P}}$ in (1) contains the query and training examples from the same training task $f$ in the set of training tasks $\mathcal{T}$. Specifically, each training task $f$ is a composition of $K$ functions $f=f_{K}\circ\cdots\circ f_{2}\circ f_{1}$ where each function $f_{k}$ belongs to a function set $\mathcal{F}$. The $k$-th step label of the query is ${\bm{z}}_{k}=f_{k}({\bm{z}}_{k-1})$ given the $k$-th step input ${\bm{z}}_{k-1}$ with ${\bm{z}}_{k}\in\mathcal{M}$, $k\in[K]$. Moreover, the $k$-th step label of the $i$-th ($i\in[l_{tr}]$) context example is ${\bm{y}}_{i,k}=f_{k}({\bm{y}}_{i,k-1})$ given the $k-1$th step input ${\bm{y}}_{i,k-1},k\in[K]$ with ${\bm{x}}_{i},{\bm{y}}_{i,k}\in\mathcal{M}$, where ${\bm{y}}_{i,0}:={\bm{x}}_{i}$ [^2] The formulation of $f$ is motivated by recent theoretical works on model training or ICL with Transformers. Please see Appendix A.2 for details..We assume that $f_{k}({\bm{x}})\neq f_{k^{\prime}}({\bm{x}}^{\prime})$ if and only if either ${\bm{x}}\neq{\bm{x}}^{\prime}$ or $f_{k}\neq f_{k^{\prime}}$.
@@
Consider a simple two-step inference example with $K=2$, ${\bm{\mu}}_{1}^{\prime}$, ${\bm{\mu}}_{2}^{\prime}$ as the TSR pattern, and ${\bm{\delta}}=0$ in inputs and outputs of every step, as shown in Figure 1. The black solid arrows denote the correct inference process, where $f_{1}({\bm{\mu}}_{1}^{\prime})={\bm{\mu}}_{1}^{\prime}$, $f_{1}({\bm{\mu}}_{2}^{\prime})={\bm{\mu}}_{2}^{\prime}$, $f_{2}({\bm{\mu}}_{1}^{\prime})={\bm{\mu}}_{2}^{\prime}$, and $f_{1}({\bm{\mu}}_{2}^{\prime})={\bm{\mu}}_{1}^{\prime}$. Hence, ${\bm{\mu}}_{1}^{\prime}\rightarrow{\bm{\mu}}_{1}^{\prime}\rightarrow{\bm{\mu}}_{2}^{\prime}$ and ${\bm{\mu}}_{2}^{\prime}\rightarrow{\bm{\mu}}_{2}^{\prime}\rightarrow{\bm{\mu}}_{1}^{\prime}$ are two inference sequences under the function $f$. The testing examples contain errors and follow the transition matrices ${\bm{A}}_{1}^{f}$ and ${\bm{A}}_{2}^{f}$ (brown dashed arrows). We let ${\bm{A}}_{1}^{f}=\begin{pmatrix}0.6&0.4\\ 0.4&0.6\end{pmatrix}$, ${\bm{A}}_{2}^{f}=\begin{pmatrix}0.4&0.6\\ 0.8&0.2\end{pmatrix}$, which results in ${\bm{B}}^{f}=\begin{pmatrix}0.56&0.44\\ 0.64&0.36\end{pmatrix}$.

### v3

@@ 1 Introduction
Transformer-based large-scale foundation models, such as GPT-3 (Brown et al., 2020), GPT-4 (OpenAI, 2023), LLaMa (Touvron et al., 2023a; b), and Sora (Liu et al., 2024), have demonstrated remarkable success across various tasks, including natural language processing (Brown et al., 2020; Touvron et al., 2023b), multimodal learning (OpenAI, 2023; Radford et al., 2021), and image/video generation (OpenAI, 2023; Liu et al., 2024). What is more surprising is that large language models (LLMs) demonstrate reasoning ability through the so-called “Chain-of-Thought” (CoT) method (Wei et al., 2022). The objective is to let a pre-trained LLM generate $K$ steps of reasoning given input query ${\bm{x}}_{query}$ without any fine-tuning. To achieve that, the input ${\bm{x}}_{query}$ is augumented with $l$ examples $\{{\bm{x}}_{i},\{{\bm{y}}_{i,j}\}_{j=1}^{K}\}_{i=1}^{l}$ of a certain $K$-step reasoning task, where each ${\bm{x}}_{i}$ is the input with ${\bm{y}}_{i,j}$ as the $j$-th reasoning step, and ${\bm{y}}_{i,K}$ is the final output. A pre-trained model then takes the resulting augmented input, referred to as a prompt, and outputs the corresponding reasoning steps $\{{\bm{z}}_{j}\}_{j=1}^{K}$ for ${\bm{x}}_{query}$, or simply outputs ${\bm{z}}_{K}$. CoT can be viewed as an extended and more intelligent method than the previous in-context learning (ICL) method, where only input-label pairs $\{{\bm{x}}_{i},{\bm{y}}_{i,K}\}_{i=1}^{l}$ are augmented in the prompt to predict ${\bm{z}}_{K}$ with the pre-trained model.
Inspired by the outstanding empirical performance of CoT in arithmetic reasoning (Wang et al., 2023; Zhang et al., 2023b; Wang & Zhou, 2024), symbolic reasoning (Zhang et al., 2023b; Zhou et al., 2023), and commonsense reasoning (Wang et al., 2023; Wang & Zhou, 2024), there have been some recent works (Li et al., 2023c; Feng et al., 2023; Li et al., 2024d; Yang et al., 2024; Wen et al., 2024) on the theoretical understanding of CoT. These works investigate CoT from the perspective of expressive power, i.e., they construct the Transformer architecture that is proven to have the CoT ability. They also demonstrate empirically that supervised training on pairs of CoT prompts and corresponding outputs can lead to models with CoT ability. However, none of these results theoretically address the question of why a Transformer can obtain generalization-guaranteed CoT ability by training from data with gradient-based methods. Meanwhile, another line of research (Zhang et al., 2023a; Huang et al., 2023; Wu et al., 2023; Li et al., 2024a) aims to unveil the reasons behind the ICL ability of Transformers through characterizing the training dynamics of a Transformer in the supervised setting. These analyses are specifically applicable to ICL. Therefore, a theoretical question still remains less explored, i.e.,
@@ 1.1 Major Contributions
Following Li et al. (2023c); Feng et al. (2023); Li et al. (2024d); Yang et al. (2024); Wen et al. (2024), we train the model in a supervised setting using prompt and label pairs. This paper provides the first theoretical analysis of the training dynamics of nonlinear Transformers to achieve CoT ability. We prove that the learned model has guaranteed CoT ability for new tasks with distribution shifts from the training tasks, even when there exist noisy and erroneous context examples in the prompt. We theoretically characterize the required number of training samples and iterations needed to train a desirable model and the number of context examples required for successful CoT reasoning with a generalization guarantee. Moreover, we provide a theoretical explanation for why CoT outperforms ICL in some cases. Our main technical contributions are as follows:
@@ 1.2 Related works
Expressive power of CoT Li et al. (2023c) proves the existence of a Transformer that can learn a multi-layer perceptron (MLP). They interpret CoT as first filtering important tokens and then making predictions by ICL. They also establish the required number of context examples for a desired prediction with the constructed Transformer. Feng et al. (2023); Li et al. (2024d); Merrill & Sabharwal (2024) show that Transformers with CoT are more expressive than Transformers without CoT. Yang et al. (2024); Wen et al. (2024) show the superiority of standard Transformers in some reasoning tasks compared with recurrent neural networks and linear Transformers.
@@
Training and Generalization of Transformers There have been several recent works about the optimization and generalization analysis of Transformers. Jelassi et al. (2022); Li et al. (2023d); Oymak et al. (2023); Li et al. (2023a; b; 2024b); Luo (2023); Huang et al. (2024); Zhang et al. (2024) study the generalization of one-layer Transformers by assuming spatial association, semantic/contextual structure, or the majority voting of tokens in the data. Oymak et al. (2023); Tarzanagh et al. (2023b; a); Tian et al. (2023a; b); Li et al. (2024c); Ildiz et al. (2024); Nichani et al. (2024); Makkuva et al. (2024b) investigate the training dynamics or loss landscape of Transformers for the next token prediction by assuming infinitely long input sequences, causal structure/Markov Chain of data, or a proper prediction head. Deora et al. (2023); Chen & Li (2024) analyze the optimization and generalization of multi-head attention networks.
@@ 2.1 Training to acquire the Chain-of-Thought ability
The learning model is a single-head, one-layer attention-only Transformer. We consider positional encoding $\{{\bm{c}}_{k}\}_{k=1}^{K}\in\mathbb{R}^{2d_{\mathcal{X}}}$. Following theoretical works (Jelassi et al., 2022; Huang et al., 2024; Ildiz et al., 2024), we add the positional encoding to each ${\bm{p}}_{i}$ by $\tilde{{\bm{p}}}_{i}={\bm{p}}_{i}+{\bm{c}}_{(i\mod K)}$, $i\in[K(l_{tr}+1)]$. $\tilde{{\bm{p}}}_{query}$ is also defined by adding the corresponding ${\bm{c}}_{k}$ to ${\bm{p}}_{query}$. Mathematically, given a prompt ${\bm{P}}$ defined in (1) with $\text{len}({\bm{P}})$ (which is at most $K(l_{tr}+1)$) denoting the number of columns, it can be written as
$F(\Psi;{\bm{P}})=\sum_{i=1}^{\text{len}({\bm{P}})-1}{\bm{W}}_{V}\tilde{{\bm{p}}}_{i}\cdot\text{softmax}(({\bm{W}}_{K}\tilde{{\bm{p}}}_{i})^{\top}{\bm{W}}_{Q}\tilde{{\bm{p}}}_{query}),$ (2)
where ${\bm{W}}_{Q},{\bm{W}}_{K}\in\mathbb{R}^{m\times(2d_{\mathcal{X}})}$, ${\bm{W}}_{V}\in\mathbb{R}^{d_{\mathcal{X}}\times(2d_{\mathcal{X}})}$ are the embedding matrices for queries, keys, and values, respectively. $\Psi:=\{{\bm{W}}_{Q}$, ${\bm{W}}_{K},{\bm{W}}_{V}\}$ is the set of all model weights [^1] We focus on a one-layer single-head Transformer motivated by recent advancements and current state in Transformer and CoT analysis. Please see Appendix B.1 for discussion. . Typically, $m>2d_{\mathcal{X}}$. Here, $\text{softmax}(({\bm{W}}_{K}\tilde{{\bm{p}}}_{i})^{\top}{\bm{W}}_{Q}\tilde{{\bm{p}}}_{query})=e^{({\bm{W}}_{K}\tilde{{\bm{p}}}_{i})^{\top}{\bm{W}}_{Q}\tilde{{\bm{p}}}_{query}}/\sum_{j=1}^{\text{len}({\bm{P}})-1}e^{({\bm{W}}_{K}\tilde{{\bm{p}}}_{j})^{\top}{\bm{W}}_{Q}\tilde{{\bm{p}}}_{query}}$.
@@ 2.2 Training Algorithm
For simplicity of analysis, we let ${\bm{W}}={\bm{W}}_{K}^{\top}{\bm{W}}_{Q}$ and ${\bm{W}}_{V}=(\bm{0}_{d_{\mathcal{X}}\times d_{\mathcal{X}}}\ {\bm{I}}_{d_{\mathcal{X}}})\in\mathbb{R}^{d_{\mathcal{X}}\times(2d_{\mathcal{X}})}$ as (Jelassi et al., 2022; Huang et al., 2023; Zhang et al., 2023a; Huang et al., 2024). Let $\{{\bm{c}}_{k}\}_{k=1}^{K}$ be a set of orthonormal vectors. The model is trained using stochastic gradient descent (SGD) with step size $\eta$ with batch size $B$, summarized in Algorithm 1 in Appendix C. Each entry of ${\bm{W}}^{(0)}$ is generated from $\mathcal{N}(0,\xi^{2})$ for a tiny $\xi>0$. ${\bm{W}}_{V}$ is fixed during the training. The fraction of prompts with ${\bm{z}}_{k-1}$ as the query input is $1/K$ by uniform sampling for any $k\in[K]$ in each batch.
@@ 2.3 Chain-of-Thought Inference
We follow the CoT-I/O scheme formulated in (Li et al., 2023c; Feng et al., 2023; Li et al., 2024d; Yang et al., 2024; Park et al., 2024) as the inference method. Specifically, for a $K$-step CoT with $l_{ts}$ examples on a certain $f\in\mathcal{T}^{\prime}$, given the testing prompt ${\bm{P}}$ defined in (4), let ${\bm{P}}_{1}={\bm{P}}$ and ${\bm{P}}_{0}$ be the first $K\cdot l_{ts}$ columns of ${\bm{P}}$. When we use CoT prompting for prediction in the $k$-th step, we first generate the output ${\bm{v}}_{k}$, $k\in[K]$ via greedy decoding by feeding the $k$-th step prompt ${\bm{P}}_{k}$ to the trained model $\Psi$ obtained from (3). The greedy decoding scheme means outputting the most probable token from the discrete set $\mathcal{Y}$ of all possible outputs, as stated in (5).
@@
where ${\bm{q}}_{k}$ is the $k$-th step reasoning column for the query. The model finally outputs ${\bm{v}}_{1},\cdots,{\bm{v}}_{K}$ as CoT result for query ${\bm{x}}_{query}$ by (5). The CoT process is summarized in Algorithm 2 of Appendix C.
When $K\geq 2$, following (Li et al., 2023c; Feng et al., 2023; Li et al., 2024d; Yang et al., 2024), the CoT generalization error given the testing query ${\bm{x}}_{query}$, the testing data distribution $\mathcal{D}^{\prime}$, and the labels $\{{\bm{z}}_{k}\}_{k=1}^{K}$ on a $K$-steps testing task $f\in\mathcal{T}^{\prime}$ is defined as
@@ 3.1 Main Theoretical Insights
We consider the setup that the model is trained using samples generated from tasks in $\mathcal{T}$ that operate on $M$ orthonormal training-relevant (TRR) patterns, while both CoT and ICL are evaluated on tasks in $\mathcal{T}^{\prime}$ that operate on $M^{\prime}$ orthonormal testing-relevant (TSR) patterns.We consider the general setup that the context examples in the prompt for CoT and ICL testing are both noisy, i.e., TSR patterns with additive noise, and partially inaccurate, i.e., the reasoning in some examples contains incorrect steps. Our main insights are as follows.
@@
P2. Guaranteed CoT Generalization. To achieve zero CoT error on tasks in $\mathcal{T}^{\prime}$ and data based on TSR patterns that contain a non-trivial component in the span of TRR patterns with the learned model, Theorem 2 shows that the required number of context examples, where noise and errors are present, for task $f$ in the testing prompt is proportional to ${(\alpha^{\prime}\tau^{f}\rho^{f})}^{-2}$. Here, $\alpha^{\prime}$ is the fraction of context examples with inputs that share the same TSR patterns as the query. $\tau^{f}$ in $(0,1)$ measures the fraction of accurate context examples, and a larger constant $\rho^{f}$ in $(0,1)$ reflects a higher reasoning accuracy in each step of the examples. This result formally characterizes the intuition that more accurate context examples and more similar examples to the query improve the CoT accuracy.
@@ 3.2 The Formulation of Data and Tasks
Every training prompt ${\bm{P}}$ in (1) contains the query and training examples from the same training task $f$ in the set of training tasks $\mathcal{T}$. Specifically, each training task $f$ is a composition of $K$ functions $f=f_{K}\circ\cdots\circ f_{2}\circ f_{1}$ where each function $f_{k}$ belongs to a function set $\mathcal{F}$. The $k$-th step label of the query is ${\bm{z}}_{k}=f_{k}({\bm{z}}_{k-1})$ given the $k$-th step input ${\bm{z}}_{k-1}$ with ${\bm{z}}_{k}\in\mathcal{M}$, $k\in[K]$. Moreover, the $k$-th step label of the $i$-th ($i\in[l_{tr}]$) context example is ${\bm{y}}_{i,k}=f_{k}({\bm{y}}_{i,k-1})$ given the $k-1$th step input ${\bm{y}}_{i,k-1},k\in[K]$ with ${\bm{x}}_{i},{\bm{y}}_{i,k}\in\mathcal{M}$, where ${\bm{y}}_{i,0}:={\bm{x}}_{i}$ [^2] The formulation of $f$ is motivated by recent theoretical works on model training or ICL with Transformers. Please see Appendix B.2 for details..We assume that $f_{k}({\bm{x}})\neq f_{k^{\prime}}({\bm{x}}^{\prime})$ if and only if either ${\bm{x}}\neq{\bm{x}}^{\prime}$ or $f_{k}\neq f_{k^{\prime}}$.
@@
Consider a simple two-step inference example with $K=2$, ${\bm{\mu}}_{1}^{\prime}$, ${\bm{\mu}}_{2}^{\prime}$ as the TSR pattern, and ${\bm{\delta}}=0$ in inputs and outputs of every step, as shown in Figure 1. The black solid arrows denote the correct inference process, where $f_{1}({\bm{\mu}}_{1}^{\prime})={\bm{\mu}}_{1}^{\prime}$, $f_{1}({\bm{\mu}}_{2}^{\prime})={\bm{\mu}}_{2}^{\prime}$, $f_{2}({\bm{\mu}}_{1}^{\prime})={\bm{\mu}}_{2}^{\prime}$, and $f_{2}({\bm{\mu}}_{2}^{\prime})={\bm{\mu}}_{1}^{\prime}$. Hence, ${\bm{\mu}}_{1}^{\prime}\rightarrow{\bm{\mu}}_{1}^{\prime}\rightarrow{\bm{\mu}}_{2}^{\prime}$ and ${\bm{\mu}}_{2}^{\prime}\rightarrow{\bm{\mu}}_{2}^{\prime}\rightarrow{\bm{\mu}}_{1}^{\prime}$ are two inference trajectories under the function $f$. The testing examples contain errors and follow the transition matrices ${\bm{A}}_{1}^{f}$ and ${\bm{A}}_{2}^{f}$ (brown dashed arrows). We let ${\bm{A}}_{1}^{f}=\begin{pmatrix}0.6&0.4\\ 0.4&0.6\end{pmatrix}$, ${\bm{A}}_{2}^{f}=\begin{pmatrix}0.4&0.6\\ 0.8&0.2\end{pmatrix}$, which results in ${\bm{B}}^{f}=\begin{pmatrix}0.56&0.44\\ 0.64&0.36\end{pmatrix}$.